OpenStudy (anonymous):

Write the equation for this situation: The cost, T, in hundreds of dollars, at one community college is 5 more than 1/2 the number of credits, c. Then graph the equation and use the graph to determine the cost of tuition if a student registers for 8 credits.

OpenStudy (phi):

can you write the equation?

OpenStudy (anonymous):

no

OpenStudy (phi):

the cost of the college is given a "name" what name did they give it?

OpenStudy (anonymous):

i dont see it giving it a name

OpenStudy (phi):

The cost, T, in hundreds of dollars,

OpenStudy (anonymous):

your saying the name is T

OpenStudy (phi):

yes. so as you read the problem The cost, T, in hundreds of dollars, at one community college is we abbreviate this to T is (throw out all the other words) in math, the word "is" often means = for example x is 10 means x=10 so T is becomes T= the start of an equation!

OpenStudy (phi):

T is 5 more than becomes T= 5+ (5 more than means we add 5 more to something)

OpenStudy (phi):

1/2 the number of credits, c. they give a "name" for the number of credits. do you see what it is?

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

c=8 right

OpenStudy (phi):

just c. so the equation in words is T is 5 more than 1/2 C can you change that into math?

OpenStudy (anonymous):

t=5+1/2c

OpenStudy (phi):

T= 5+ c/2 or T= C/2 +5 that is the equation of a line (but c for x, and T for y, but that's ok) Then graph the equation I would label the x-axis C (credits), and the vertical axis T (cost) can you plot the line?

OpenStudy (phi):

yes 1/2 times c plus 5 now you must plot it. one easy point is when C=0: t= 5+(1/2)*0= 5 so (0,5) is one point.

OpenStudy (phi):

another point might be when C=2 T= (1/2)*2+5 = 1+5= 6 so (2,6) is another point.

OpenStudy (anonymous):

ok where does the 8 can in

OpenStudy (anonymous):

Hi {{T[H] = (1/2)c + 5 8 credits, then cost is $900.

OpenStudy (phi):

use the graph to determine the cost of tuition if a student registers for 8 credits. they want you to draw the line, then find where C=8 on the line and read off the T value.

OpenStudy (anonymous):

@amber517 hope this works =)

OpenStudy (anonymous):

huh

OpenStudy (phi):

can you graph the line t=5+(1/2)*c

OpenStudy (anonymous):

idk

OpenStudy (phi):

if it were y= (1/2)*x + 5 could you graph it?

OpenStudy (anonymous):

ya i thinkso

OpenStudy (phi):

after plotting y= (1/2)*x + 5 rename the x axis C rename the y axis T (because the problem uses those names rather than x and y)

OpenStudy (anonymous):

y=1/2 x=5

OpenStudy (phi):

no, when you have y= (1/2)*x + 5 you can replace the x with a number (try 0) to get y (see the problem you worked on before this). if x=0 can you find y? you are trying to find (0, ) (fill in the blank) then plot that point. then try x=2, find y and plot that point. connect the dots to get a line

OpenStudy (anonymous):

0,5 2,6

OpenStudy (phi):

yes plot those points and draw a line through them. then label the x-axis C and the y-axis T (to match the problem)

OpenStudy (anonymous):

ok but what isthe 8 for not what i' understandng

OpenStudy (phi):

First, did you plot the line?

OpenStudy (phi):

Look at justin's graph

OpenStudy (anonymous):

ok

OpenStudy (phi):

after you plot the line, find 8 on the C axis (x axis), and follow it up until you reach the line. put a dot there on the line. then from the dot, move sideways to the left until you reach the T axis (y axis). read the number on the T axis.

OpenStudy (anonymous):

7

OpenStudy (phi):

 are you sure? I got 9